OpenStudy (anonymous):

Solve using the principle of zero products. (11s+6)(5s-15)=0

jimthompson5910 (jim_thompson5910):

(11s+6)(5s-15)=0 11s+6=0 or 5s-15=0 11s=-6 or 5s=15 s=-6/11 or s=15/5 s=-6/11 or s=3